# KompeteAI: Accelerated Autonomous Multi-Agent System for End-to-End Pipeline Generation for Machine Learning Problems

### v1

@@ Abstract
Recent Large Language Model (LLM)-based AutoML systems demonstrate impressive capabilities but face significant limitations such as constrained exploration strategies and a severe execution bottleneck. Exploration is hindered by one-shot methods lacking diversity and Monte Carlo Tree Search (MCTS) approaches that fail to recombine strong partial solutions. The execution bottleneck arises from lengthy code validation cycles that stifle iterative refinement. To overcome these challenges, we introduce KompeteAI, a novel AutoML framework with dynamic solution space exploration. Unlike previous MCTS methods that treat ideas in isolation, KompeteAI introduces a merging stage that composes top candidates. We further expand the hypothesis space by integrating Retrieval-Augmented Generation (RAG), sourcing ideas from Kaggle notebooks and arXiv papers to incorporate real-world strategies. KompeteAI also addresses the execution bottleneck via a predictive scoring model and an accelerated debugging method, assessing solution potential using early stage metrics to avoid costly full-code execution. This approach accelerates pipeline evaluation 6.9 times. KompeteAI outperforms leading methods (e.g., RD-agent, AIDE, and Ml-Master) by an average of 3% on the primary AutoML benchmark, MLE-Bench. Additionally, we propose Kompete-bench to address limitations in MLE-Bench, where KompeteAI also achieves state-of-the-art results.
@@ 1 Introduction
Recent research has shifted towards the use of Large Language Models (LLM) as the core reasoning engine for AutoML frameworks, enabling the autonomous generation and testing of end-to-end pipelines that adapt to specific tasks (Li et al. 2024; Jiang et al. 2025; Chi et al. 2024; Liu et al. 2025; Trirat, Jeong, and Hwang 2024; Grosnit et al. 2024; Yang et al. 2025). However, these approaches have critical limitations. Initial ”one-shot” generation can yield diverse ideas but lacks iterative refinement, so a single flawed component - like suboptimal feature engineering — can undermine the entire pipeline without a way to correct it.
@@
More adaptive frameworks based on Monte Carlo Tree Search (MCTS) address this by exploring a tree of potential solutions, but often use constrained exploration and struggle to recombine promising ideas from different branches. Even recent frameworks that combine exploratory search with LLM-based reasoning (Liu et al. 2025), despite achieving state-of-the-art results on MLE-Bench (Chan et al. 2024), are limited by architectural constraints. They may not systematically preserve or merge valuable insights from distant high-performing branches. As a result, valuable solution components can be prematurely discarded.
@@
Even an ideal internal recombination can only reshuffles known ideas. Retrieval‑Augmented Generation (RAG) breaks this limit by injecting external, domain-specific knowledge — enabling the agent to explore solutions beyond its pretrained hypothesis space. This potential remains largely untapped in current implementations. Most systems apply RAG only during early stages; as the problem shifts — from statistical feature engineering to physics-based simulation or chemistry-driven descriptors — the agent cannot retrieve fresh, relevant knowledge, leading to knowledge decay as the pipeline evolves.
@@
Beyond search and knowledge limitations, all frameworks suffer from a severe execution bottleneck. Validating a single solution requires full code execution, which can take hours for complex models. This issue becomes even more problematic during debugging, as errors late in the pipeline force a complete retraining for every fix. As a result, the slow feedback loop discourages major changes and presents a serious scalability challenge.
@@
In this work, we introduce KompeteAI, an autonomous multi-agent framework for structured, multistage pipeline generation. The key innovations compared to prior approaches is presented in Table 1. To efficiently explore and exploit the search space, we employ two core operators. Adding, which dynamically generates novel stage-specific ideas by querying external knowledge sources via an adaptive RAG module, and merging, which intelligently combines the most successful solutions. We address execution bottlenecks through a predictive scoring model that prunes weak solutions early, alongside an accelerated debugging paradigm using simplified code and smaller data samples, dramatically shortening the feedback loop.
@@
Our experiments show that KompeteAI sets a new state-of-the-art on MLE-Bench, outperforming prior methods by an average of 3%, while accelerating the average test-time performance by a factor of 6.9.
@@
  • Stage-Decomposed Multi-Agent Architecture: A state-of-the-art framework that partitions the ML workflow into discrete stages, enabling agents to specialize in focused tasks, dynamically integrate external knowledge sources to enhance exploration diversity, and systematically recombine optimal partial solutions through novel addition and merging operators.
@@ 2 Related Work
Classic AutoML
Classic AutoML frameworks - TPOT, AutoGluon, AutoKeras, LightAutoML, and others (Olson and Moore 2016; Erickson et al. 2020; Jin, Song, and Hu 2019; Vakhrushev et al. 2021; Feurer et al. 2022; LeDell and Poirier 2020; Thornton et al. 2013) — automate data preprocessing, model selection, and hyperparameter tuning via heuristic search, ensembling, and Bayesian optimization. Despite their effectiveness, these systems operate within static search spaces, require manual data preparation and adaptation for each new task, and lack the dynamic coordination and continual learning capabilities inherent to multi‑agent AutoML architectures.
AutoML based LLM
LLM-based AutoML systems have progressed rapidly, offering increasingly autonomous capabilities through dynamic coordination and iterative planning (Li et al. 2024; Chi et al. 2024; Jiang et al. 2025; Trirat, Jeong, and Hwang 2024; Yang et al. 2025; Liu et al. 2025). To analyze and compare these systems meaningfully, we focus on four key aspects that consistently shape their performance and flexibility: exploration strategy, RAG, merge methods, and debugging techniques. These components address critical challenges related to how systems explore the design space, incorporate external knowledge, manage idea diversity, and implement solutions both efficiently and reliably. Notably, RD-Agent’s merge approach amounts to simple, uncontrolled recombination driven by LLM rather than a structured merge algorithm, which results in incoherent or suboptimal pipeline integrations. A summary of how existing systems approach these dimensions is provided in Table 1.
Table 1: Comparison of LLM‑based AutoML systems. Phase‑wise chain: sequential, stage‑by‑stage pipeline construction; without context: searches that ignore information from parallel or prior branches; Multi‑stage expansion: KompeteAI’s staged node growth at every pipeline level; R&D‑phase: retrieval only during initial planning, versus Dynamic RAG: on‑the‑fly literature/web queries; Incremental debugging: iterative draft–debug–refine cycles, versus Multistage debugging: KompeteAI’s multi‑agent, stage‑based error checking; Simple recombination: implicit LLM‑driven fusion of ideas, versus Controlled merger: KompeteAI’s explicit algorithmic branch combination.
Method Exploration Strategy RAG system Debugging method Merge Method
AutoKaggle Phase-wise chain Three-stage phase & Unit tests
SELA MCTS (UCT) without context
AIDE MCTS (UCT) without context Incremental
@@
R&D-Agent Multi-trace exploration R&D-phase Incremental Simple recombination
ML-Master Controlled MCTS with memory Incremental
KompeteAI Multi-stage expansion Dynamic Multistage Controlled merger
Scoring Model
Our approach to model scoring is inspired by similar performance prediction methods developed in Neural Architecture Search (Elsken, Metzen, and Hutter 2019). In NAS, a common approach is to use weight-sharing supernets, where multiple architectures are jointly trained by sharing parameters with a large model, and performance is estimated by evaluating sampled architectures on a validation set (Jawahar et al. 2023a). This method, however, faces challenges such as weight co-adaptation (Bender et al. 2018), capacity bottlenecks (Jawahar et al. 2023b), and gradient conflicts (Gong and Wang 2022). More recently (Jawahar et al. 2023a), demonstrated that LLMs can serve as effective performance predictors in NAS tasks, providing a promising alternative to traditional methods. In contrast to NAS, the AutoML setting typically involves a much broader and less constrained search space, which motivates the exploration of LLM-based performance prediction beyond neural architectures.
Benchmarks
Using Kaggle competitions to evaluate autonomous ML systems has become a modern approach in a number of recent benchmarks due to clear metrics, variety of tasks and the ability to compare with human solutions. One of the first benchmarks to systematically evaluate autonomous ML agents on Kaggle competitions was MLAgentBench (Huang et al. 2023), which focused on a small set of tasks with simple baselines, measuring agents’ ability to improve upon them. Later, DSBench (Jing et al. 2024), expanded the scope but often relied on automated filtering, which excluded many complex or non-standard competitions. The most recent effort, MLE-Bench (Chan et al. 2024) stands out for its scale and diversity, present a more challenging and realistic testbed for multi-agent AutoML systems. However, MLE-Bench also faces notable limitations, including its large size (3.3 TB) and the fact that it constructs its test sets by partitioning the original training data and then compares the resulting scores to the private leaderboard positions that cause evaluation bias.
@@ 3 KompeteAI
The pipeline of KompeteAI is demonstrated in Figure 1. It is designed to ensure robust, leak-free data handling, efficient exploration of modeling ideas, and rapid iteration, all while maintaining high code quality and logical consistency. This section outlines the key stages and mechanisms that underpin the operation of our system. It consists of four main components: Pipeline Setup, The Ideation Process, Tree-Guided Exploration, which includes operations such as node addition and merging, and the Scoring Model, designed to accelerate overall pipeline evaluation.
In our framework, each node represents a segment of the final code, and each tree level corresponds to a distinct pipeline component: Exploratory Data Analysis (EDA), Feature Engineering (FE), or Model Training (MT). Connections between nodes indicate that they form parts of the same pipeline. The main stage unfolds in three phases: tree initialization, adding, and merging. Tree initialization constructs the primary tree structure. Subsequently, within the allotted time budget, the system alternates between the adding phase — where new ideas are injected into the most promising branches — and the merging phase, which combines two promising ideas into a single, more powerful solution. The EDA phase is represented by a single root node, which aggregates all insights derived from data exploration, such as visualizations, distribution analyses, and summary statistics. This EDA node is not static: as the search progresses, it can be dynamically enriched with additional analyses prompted by downstream discoveries. In contrast, nodes at the FE and MT levels represent concrete instantiations of their respective phases, with edges between nodes indicating their inclusion in the same candidate solution. While a node may have multiple children, it is restricted to a single parent.
@@ 3.1 Pipeline Setup
This phase sets up the core components required for the next stages. The dataset is ingested by the Reader Agent, which analyzes its structure, produces a detailed task specification, and initializes the data for the RAG based on this description. The Metric Agent constructs unit tests to support submission validation and defines the evaluation metric function. The Validator Agent partitions the data according to the task specification and applies appropriate preprocessing methods to ensure a valid and reliable evaluation protocol. The Baseliner Agent generates an initial solution, establishing a lower-bound reference for expected performance. Based on the baseline score, it also assesses the quality of the data split and, if necessary, can trigger the Validator Agent to re-partition the data using an alternative strategy.
3.2 The Ideation Process
@@
Figure 1: The KompeteAI AutoML pipeline consists of three main stages: Pipeline Setup, Tree Initialization, and the Main Loop. Pipeline Setup involves data ingestion, initial analysis, metric evaluation, data validation, and establishing performance benchmarks. Tree Initialization uses EDA to generate initial insights, forming a tree structure where Feature Engineering nodes are parents to Model Training nodes. The Main Loop iteratively refines the solution tree through adding new nodes, with their performance subsequently evaluated via code acceleration, debugging, and return speed to get a score; backpropagation for scoring FE ideas based on downstream model feedback; and merging strong partial solutions to recombine effective strategies
Insighter
@@
  • Tree Memory: The agent employs a memory module over the entire ideation tree, analyzing previously generated nodes using diversity-driven strategies based on embedding similarity — such as selecting nodes closest or farthest from the parent, or sampling randomly. A top-$n$ subset is selected for inclusion in the context, and the memory is continuously updated as new nodes are added, enabling dynamic promotion of diversity.
@@ 3.3 Tree-Guided Exploration
This section provides an explanation of the core principles underlying operations on the ideation tree.
Tree Initialization
The primary objective of the tree initialization stage is to generate an initial set of candidate pipelines, providing a diverse foundation from which further exploration can proceed. This stage seeds the search space with promising ideas and establishes the initial structure for subsequent optimization. After the initial tree is constructed, we perform a backpropagation step to propagate performance signals from the model training nodes (where direct evaluation scores are available) up through the tree, updating each node’s average score based on new information to improve their representativeness within the search tree.
Adding
@@
The entire process is conditioned on a global context representation $\mathbf{c}_{t}$, which captures the current pipeline state and external knowledge in a structured and denoised form. We define the context vector as:
$\mathbf{c}_{t}=\phi_{\mathrm{EDA}}(T_{t})+\phi_{\mathrm{reader}}(T_{t})+\phi_{\mathrm{ext}}(\texttt{QueryExternal}())$
where each mapping $\phi:\mathcal{X}\rightarrow\mathcal{C}$ transforms raw, unstructured, or noisy inputs into structured semantic representations in a common context space $\mathcal{C}$. Specifically:
  • $\phi_{\mathrm{EDA}}$ encodes statistical and structural information extracted from exploratory data analysis;
  • $\phi_{\mathrm{reader}}$ derives contextual insights from competition metadata and prior solutions;
  • $\phi_{\mathrm{ext}}$ incorporates knowledge retrieved from external sources (e.g., Kaggle, arXiv).
Based on the aggregated context $\mathbf{c}_{t}$, the agent performs a structured expansion of the ideation tree in three successive steps. First, a set of candidate FE nodes is sampled from the conditional distribution
$q_{\mathrm{FE}}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{FE}}),\quad q_{\mathrm{FE}}(\cdot\mid\mathbf{c}_{t})$
where $\mathcal{V}_{\mathrm{FE}}$ denotes the space of available FE transformations. Then, for each sampled FE node $v^{\mathrm{FE}}\in\mathcal{V}_{\mathrm{FE}}$, the agent generates a corresponding set of MT nodes by sampling from
$q_{\mathrm{MT}}:\mathcal{V}_{\mathrm{FE}}\rightarrow\mathcal{P}(\mathcal{V}_{\mathrm{MT}})$
which specifies a distribution over the MT configuration space $\mathcal{V}_{\mathrm{MT}}$. Finally, a subset of FE nodes from the current tree (including both existing and newly added ones) is selected based on their scores, reflecting their estimated utility. For each selected FE node, additional MT nodes are appended as children.
The full procedure is detailed in Appendix A1.
Merging

### v2

@@ Abstract
Recent Large Language Model (LLM)-based AutoML systems demonstrate impressive capabilities but face significant limitations such as constrained exploration strategies and a severe execution bottleneck. Exploration is hindered by one-shot methods lacking diversity and Monte Carlo Tree Search (MCTS) approaches that fail to recombine strong partial solutions. The execution bottleneck arises from lengthy code validation cycles that stifle iterative refinement. To overcome these challenges, we introduce KompeteAI, a novel AutoML framework with dynamic solution space exploration. Unlike previous MCTS methods that treat ideas in isolation, KompeteAI introduces a merging stage that composes top candidates. We further expand the hypothesis space by integrating Retrieval-Augmented Generation (RAG), sourcing ideas from Kaggle notebooks and arXiv papers to incorporate real-world strategies. KompeteAI also addresses the execution bottleneck via a predictive scoring model and an accelerated debugging method, assessing solution potential using early-stage metrics to avoid costly full-code execution. This approach accelerates pipeline evaluation 6.9 times. KompeteAI outperforms leading methods (e.g., RD-agent, AIDE, and ML-Master) by an average of 3% on the primary AutoML benchmark, MLE-Bench. Additionally, we propose Kompete-bench to address limitations in MLE-Bench, where KompeteAI also achieves state-of-the-art results.
@@ 1 Introduction
Recent research has shifted towards the use of Large Language Models (LLM) as the core reasoning engine for AutoML frameworks, enabling the autonomous generation and testing of end-to-end pipelines that adapt to specific tasks (Li et al. (2024); Jiang et al. (2025); Chi et al. (2024); Liu et al. (2025); Trirat et al. (2024); Grosnit et al. (2024); Yang et al. (2025); Nam et al. (2025)). However, these approaches have critical limitations. Initial ”one-shot” generation can yield diverse ideas but lacks iterative refinement, so a single flawed component - like suboptimal feature engineering — can undermine the entire pipeline without a way to correct it.
@@
More adaptive frameworks based on Monte Carlo Tree Search (MCTS) address this by exploring a tree of potential solutions but often use constrained exploration and struggle to recombine promising ideas from different branches. Even recent frameworks that combine exploratory search with LLM-based reasoning Liu et al. (2025), despite achieving state-of-the-art results on MLE-Bench Chan et al. (2024), are limited by architectural constraints. They may not systematically preserve or merge valuable insights from distant high-performing branches during the exploration process. As a result, valuable solution components can be prematurely discarded.
@@
Even perfect internal recombination only reshuffles known ideas. Retrieval-Augmented Generation (RAG) overcomes this by injecting external, domain-specific knowledge, enabling exploration beyond the pretrained hypothesis space. Yet current systems rarely exploit this potential: most apply RAG only in the early stages Trirat et al. (2024); Yang et al. (2025), causing retrieval to fail at providing fresh knowledge and leading to knowledge decay across the pipeline.
@@
All frameworks also face a severe execution bottleneck. Validating a single solution requires full code execution, often taking hours. Debugging worsens this: late errors force complete retraining, slowing feedback and discouraging major changes. This creates a serious scalability challenge.
@@
In this work, we introduce KompeteAI, an autonomous multi-agent framework for structured, multistage pipeline generation. The key innovations compared to prior approaches are presented in Table 1. To efficiently explore and exploit the search space, we employ two core operators. Adding, which dynamically generates novel stage-specific ideas by querying external knowledge sources via an adaptive RAG module, and merging, which intelligently combines the most successful solutions. We address execution bottlenecks through a predictive scoring model that prunes weak solutions early, alongside an accelerated debugging paradigm using simplified code and smaller data samples, dramatically shortening the feedback loop. These innovations accelerate the average test-time performance by a factor of 6.9.
@@
Our experiments show that KompeteAI sets the new state-of-the-art on MLE-Bench, outperforming prior methods by an average of 3%.
@@
  • Stage-Decomposed Multi-Agent Architecture: A state-of-the-art framework that partitions the ML workflow into discrete stages, enabling agents to specialize in focused tasks, dynamically integrate external knowledge sources to enhance exploration diversity, and systematically recombine optimal partial solutions through novel adding and merging operators.
@@ 2 Related Work
Classic AutoML. Classic AutoML frameworks - TPOT, AutoGluon, AutoKeras, LightAutoML, and others Olson & Moore (2016); Erickson et al. (2020); Jin et al. (2019); Vakhrushev et al. (2021); Feurer et al. (2022); LeDell & Poirier (2020); Thornton et al. (2013) — automate data preprocessing, model selection, and hyperparameter tuning via heuristic search, ensembling, and Bayesian optimization. Despite their effectiveness, these systems operate within static search spaces, require manual data preparation and adaptation for each new task, and lack the dynamic coordination and continual learning capabilities inherent to multi‑agent AutoML architectures.
LLM-based AutoML. LLM-based AutoML systems have progressed rapidly, offering increasingly autonomous capabilities through dynamic coordination and iterative planning (AutoKaggle Li et al. (2024), SELA Chi et al. (2024), AIDE Jiang et al. (2025), AutoML-Agent Trirat et al. (2024), RD-agentYang et al. (2025), ML-Master Liu et al. (2025), MLE-STAR Nam et al. (2025)). To analyze and compare these systems meaningfully, we focus on four key aspects. The results are presented in Table 1. Notably, RD-Agent’s merging relies on uncontrolled LLM-driven recombination, often producing incoherent or suboptimal integrations, while MLE-STAR’s single, terminal merging step simply consolidates ideas without meaningfully improving solution exploration.
Exploration RAG Debuging Merging
SELA MCTS (UCT) w/o context
AIDE MCTS (UCT) w/o context Incremental
@@
R&D-Agent Multi-trace exploration R&D-phase Incremental Recombination
ML-Master MCTS (UCT) with context Incremental
MLE-STAR Incremental exploration Single-stage Final model ensembling
[…]
Table 1: Comparison of LLM-based AutoML systems.
Scoring Model. Our approach of performance scoring is inspired by similar performance prediction methods developed in Neural Architecture Search Elsken et al. (2019). In NAS, a common approach is to use weight-sharing supernets, where multiple architectures are jointly trained by sharing parameters with a large model, and performance is estimated by evaluating sampled architectures on a validation set Jawahar et al. (2023a). This method, however, faces challenges such as weight co-adaptation Bender et al. (2018), capacity bottlenecks Jawahar et al. (2023b), and gradient conflicts Gong & Wang (2022). More recently Jawahar et al. (2023a), demonstrated that LLMs can serve as effective performance predictors in NAS tasks, providing a promising alternative to traditional methods. In contrast to NAS, the AutoML setting typically involves a much broader and less constrained search space, which motivates the exploration of LLM-based performance prediction beyond neural architectures.
Benchmarks. Using Kaggle competitions to evaluate autonomous ML systems has become a modern approach in a number of recent benchmarks due to clear metrics, variety of tasks, and the ability to compare with human solutions. One of the first benchmarks for systematically evaluating autonomous ML agents in Kaggle competitions was MLAgentBench Huang et al. (2023), which focused on a small set of tasks with simple baselines, measuring agents’ ability to improve on them. Later, DSBench Jing et al. (2024), expanded the scope but often relied on automated filtering, which excluded many complex or non-standard competitions. The most recent effort, MLE-Bench Chan et al. (2024) stands out for its scale and diversity, presenting a more challenging and realistic testbed for multi-agent AutoML systems. However, MLE-Bench also faces notable limitations, including its large size (3.3 TB) and the fact that it constructs its test sets by partitioning the original training data and then compares the resulting scores to the private leaderboard positions that cause evaluation bias.
@@ 3 KompeteAI
The pipeline of KompeteAI is demonstrated in Figure 1. It is designed to ensure robust, leak-free data handling, efficient exploration of modeling ideas, and rapid iteration, all while maintaining high code quality and logical consistency. This section outlines the key components and mechanisms that underpin the operation of our system. It consists of three main stages: (3.1) Pipeline Setup, which prepares the core components; (3.2) Tree Initialization, which generates an initial set of candidate pipelines; and (3.3) Tree-Guided Exploration, the main stage involving tree operations such as node adding and merging, supported by a Scoring Model to speed up pipeline evaluation.
@@ 3.1 Pipeline Setup
Figure 1: The KompeteAI AutoML pipeline.
This phase sets up the core components required for the next stages. The dataset is ingested by The Reader Agent, which analyzes its structure, produces a detailed task specification, and initializes the data for the RAG based on this description. The Metric Agent constructs unit tests to support submission validation and defines the evaluation metric function. The Validator Agent partitions the data according to the task specification and applies appropriate pre-processing methods to ensure a valid and reliable evaluation protocol. The Baseliner Agent generates an initial solution, establishing a lower-bound reference for expected performance. Based on the baseline score, it also assesses the quality of the data split and, if necessary, can trigger The Validator Agent to re-partition the data using an alternative strategy.
3.2 Tree Initialization
The primary objective of the tree initialization stage is to generate an initial set of candidate pipelines, providing a diverse foundation from which further exploration can proceed. This stage seeds the search space with promising ideas and establishes the initial structure for subsequent optimization.
3.2.1 Node Representation
In our framework, each node encodes a code segment, and each tree level maps to a pipeline component: Exploratory Data Analysis (EDA), Feature Engineering (FE), or Model Training (MT). Connections between nodes define candidate pipelines. The root EDA node aggregates insights from data exploration — visualizations, distributions, and summary statistics — and can be dynamically enriched as the search progresses. In contrast, FE and MT nodes represent concrete instantiations of their phases, with edges marking their inclusion in the same solution. Each node may have multiple children but only one parent.
3.2.2 The Ideation Process
@@
  • Tree Memory: The agent employs a memory module over the entire ideation tree, analyzing previously generated nodes using diversity-driven strategies based on embedding cosine similarity — such as selecting nodes closest or farthest from the parent, or sampling randomly. A top-$n$ subset is selected for inclusion in the context, and the memory is continuously updated as new nodes are added, enabling dynamic promotion of diversity.
@@ 3.3 Tree-Guided Exploration
Tree-Guided Exploration has three components: Adding, Merging, and Scoring Model. The Tree Initialization stage builds the primary structure. Within the time budget, the system alternates between Adding — injecting new ideas into promising branches — and Merging, which combines two strong ideas into a single solution. The scoring model works alongside these phases to accelerate pipeline evaluation. After each Adding–Merging cycle, we backpropagate performance signals from MT nodes (where evaluation scores are available) upward, updating each node’s average score to improve their representativeness in the search tree. The average score of each node is calculated as the mean of the scores of its children.
3.3.1 Adding
@@
The process is conditioned on a global context vector $\mathbf{c}_{t}$, representing the current pipeline state and external knowledge in structured form.
We define $\mathbf{c}_{t}=\phi_{\mathrm{EDA}}(T_{t})+\phi_{\mathrm{reader}}(T_{t})+\phi_{\mathrm{ext}}(\texttt{QueryExternal}())$, where each mapping $\phi:\mathcal{X}\rightarrow\mathcal{C}$ converts noisy inputs into structured semantics in the shared context space $\mathcal{C}$. Here, $\phi_{\mathrm{EDA}}$ encodes statistical/structural signals from exploratory analysis, $\phi_{\mathrm{reader}}$ extracts insights from metadata and past solutions, and $\phi_{\mathrm{ext}}$ adds knowledge from external sources such as arXiv or Kaggle.
Based on the aggregated context $\mathbf{c}_{t}$, the agent performs a structured expansion of the ideation tree in three successive steps. First, it samples a set of candidate FE nodes according to the context $\mathbf{c}_{t}$, where $\mathcal{V}_{\mathrm{FE}}$ denotes the space of available FE transformations. Then, for each sampled FE node $v^{\mathrm{FE}}$, the agent generates a corresponding set of MT nodes within the MT configuration space $\mathcal{V}_{\mathrm{MT}}$. Finally, a subset of FE nodes from the current tree (including both existing and newly added ones) is selected based on their scores, and for each selected node additional MT nodes are appended as children.
The full procedure is detailed in Appendix 1.
3.3.2 Merging